The market implied spot rates 3.5-years and 4-years from today are 5.00% and 5.25% respectively. Calculate the forward rate three and a half years from today and express it in annual terms.
Answer #1

Calculating rate 3.5 years from now,

(1 + r/2) = (1 + 0.0525/2)8/(1 + 0.05/2)7

r/2 = 3.50%

So,

APR = 2(0.035)

APR = 7.00%